# Hydrodynamics as the effective field theory of strong-to-weak spontaneous symmetry breaking

### v1

@@ 1 Introduction
This paper synthesizes the story of SWSSB with the EFT of hydrodynamics, focusing on the simplest case of diffusion for clarity, although our punchline should directly generalize to other hydrodynamic models. Our key observation, which we will carefully and pedagogically motivate, is that SWSSB quantitatively justifies the reparameterization symmetry of the EFT, providing a physical justification for the one “mysterious” aspect of the prescription of Crossley et al. (2017). Aspects of this idea, in particular an identification of broken strong symmetry in hydrodynamics, can be found in earlier work Akyuz et al. (2024); Ogunnaike et al. (2023). Our paper will start by deducing the most general effective field theory for “superfluidity” in a theory with SWSSB, and derive the corresponding EFT, which exactly reproduces hydrodynamics Crossley et al. (2017). Along the way, we will also propose a physical interpretation for the integrated-in $\phi$ fields as quantum channels, which may be of independent interest.
@@ 2 Review: SSB in pure states
for any $t$-independent function $f$. This enforces that $\mathcal{H}=\mathcal{H}(n)$ does not depend at all on $\phi$. The resulting Hamiltonian EFT is therefore trivially integrable: $\partial_{t}n=0$ and $\partial_{t}\phi$ is independent of $t$.
@@ 4 SWSSB in a “doubled” Hilbert space
$|a\rangle\langle b|\rightarrow|ab)$ (30)
The parentheses bra-ket notation emphasizes that we are in the doubled Hilbert space. The inner product on the doubled Hilbert space is
@@
$B_{\mathrm{R}}|ab)$ $:=|a\rangle\langle b|B^{\mathsf{T}}.$ (32b)
@@
where
@@
Define the time-reversed superoperator
@@
$\tilde{n}_{i\mathrm{a}}$ $=\tilde{n}_{i\mathrm{a}},$ (50a)
$\tilde{n}_{i\mathrm{r}}$ $=\tilde{n}_{i\mathrm{r}},$ (50b)
@@
Now, let us explain why there is no Mermin-Wagner Theorem Mermin and Wagner (1966) for SWSSB. Roughly speaking, from the eigenstate thermalization hypothesis and statistical mechanics, insofar as local correlation functions are concerned, the density matrix in our mixed steady state $\rho$ is well-approximated by a grand canonical ensemble (7), which exhibits SWSSB. Since (7) is a tensor product, there is no long-range order acccessible in any local correlators, without using multiple copies of the density matrix. (This renders the probes of SSB unmeasurable in current experiments.)
Let us now be a little more careful. To see the emergence of a “long-range order” which can be associated to SWSSB, we should start with a strongly symmetric state, and a useful choice will be (42). In the thermodynamic limit, if $N=\langle n\rangle V$ (where $V$ is the number of lattice sites) is chosen such that $\langle n\rangle$ is the expectation value in the grand canonical ensemble $\rho$, defined in (35). To see the presence of long-range order, we must look at more complicated correlation functions involving multiple copies of a density matrix, such as

### v2

@@ 1 Introduction
This paper synthesizes the story of SWSSB with the EFT of hydrodynamics, focusing on the simplest case of diffusion for clarity, although our punchline should directly generalize to other hydrodynamic models. Our key observation, which we will carefully and pedagogically motivate, is that SWSSB quantitatively justifies the reparameterization symmetry of the EFT, providing a physical justification for the one “mysterious” aspect of the prescription of Crossley et al. (2017). Aspects of this idea, in particular an identification of broken strong symmetry in hydrodynamics, can be found in earlier work Akyuz et al. (2024); Ogunnaike et al. (2023); Moudgalya and Motrunich (2023). Our paper will start by deducing the most general effective field theory for “superfluidity” in a theory with SWSSB, and derive the corresponding EFT, which exactly reproduces hydrodynamics Crossley et al. (2017). Along the way, we will also propose a physical interpretation for the integrated-in $\phi$ fields as quantum channels, which may be of independent interest.
@@ 2 Review: SSB in pure states
for any $t$-independent function $f$. This enforces that $\mathcal{H}=\mathcal{H}(n)$ does not depend at all on $\phi$, which is in turn a consequence of the fact that (outside of a “healing length”) there is no energy penalty to phase fluctuations in a normal fluid. The resulting Hamiltonian EFT is therefore trivially integrable: $\partial_{t}n=0$ and $\partial_{t}\phi$ is independent of $t$. If the system is anomalous, there is an additional term $\mathcal{L}_{\mathrm{anom}}\sim c\partial_{t}\phi\partial_{x}\phi$ in $d=1$ spatial dimension, that is allowed, as it shifts by a total derivative under (19); see Appendix A. Since we are not interested in the anomaly, we will set the coefficient $c=0$ for the remainder of the paper.
@@ 4 SWSSB in a “doubled” Hilbert space
$|a\rangle\langle b|\rightarrow|ab),$ (30)
where $|ab)\equiv|a\rangle\otimes|b\rangle^{*}$ with $|b\rangle^{*}$ being the complex conjugation of the state $|b\rangle$. The parentheses bra-ket notation emphasizes that we are in the doubled Hilbert space. The inner product on the doubled Hilbert space is
@@
$B_{\mathrm{R}}|ab)$ $:=|a\rangle\langle b|B^{\mathsf{T}},$ (32b)
@@
where the first equality comes from the identity $\mathrm{tr}(\mathcal{A}\rho)^{*}=\mathrm{tr}(\rho^{\dagger}\mathcal{A}^{\dagger})$, and in the second equality we introduced
@@
Note the transformation in (47) is sometimes called modular conjugation. Define the time-reversed superoperator
@@
$\tilde{n}_{i\mathrm{a}}$ $=n_{i\mathrm{a}},$ (50a)
$\tilde{n}_{i\mathrm{r}}$ $=n_{i\mathrm{r}},$ (50b)
@@
Now, let us explain why there is no Mermin-Wagner Theorem Mermin and Wagner (1966) for SWSSB. Roughly speaking, from the eigenstate thermalization hypothesis and statistical mechanics, insofar as local correlation functions are concerned, the density matrix in our mixed steady state $\rho$ is well-approximated by a grand canonical ensemble (7), which exhibits SWSSB. Since (7) is a tensor product, there is no long-range order accessible in any local correlators, without using multiple copies of the density matrix. (This renders the probes of SSB unmeasurable in current experiments.)
Let us now be a little more careful. To see the emergence of a “long-range order” which can be associated to SWSSB, we should start with a strongly symmetric state, and a useful choice will be (42). In the thermodynamic limit, if $N=\langle n\rangle V$ (where $V$ is the number of lattice sites) is chosen such that $\langle n\rangle$ is the expectation value in the grand canonical ensemble $\rho$, defined in (35). To see the presence of long-range order, we must look at more complicated correlation functions involving multiple copies of a density matrix, such as the Wightman’s or Rényi-1 correlator Liu et al. (2024); Weinstein (2024)
@@ Acknowledgements
Appendix A 1+1d gapless Dirac fermion
A symmetric quantum system can have nontrivial excitations if the symmetry is anomalous. For example, the Dirac fermion in 1+1 dimensions could be gapless without breaking any symmetry. This example does not appear to fall into our categories as it has neither SSB nor reparametrization symmetry. However, we show here that the EFT can be written in a way with manifest reparametrization symmetry. The key insight is that the Dirac fermion has two $U(1)$ symmetries with a mixed anomaly; treating the two $U(1)$ symmetries on equal footing leads to an action with manifest reparametrization symmetry. In the following, we will work with the $1+1d$ compact boson, which is related to the $1+1$d Dirac fermion by bosonization, and is completely equivalent at the level of hydrodynamics.
We begin with the action for the compact boson with the identification $\varphi\sim\varphi+2\pi$
$S[\varphi]=\frac{R^{2}}{4\pi}\int\mathrm{d}x\mathrm{d}t~{}(\partial_{t}\varphi)^{2}-(\partial_{x}\varphi)^{2}.$ (91)
The radius $R=\sqrt{2}$ corresponds to the bosonized free Dirac fermion. This action is manifestly invariant under the symmetry $\varphi\to\varphi+c$. It will be helpful to review how T-duality rewrites the action in terms of a dual field $\theta$ which is manifestly invariant under the symmetry $\theta\to\theta+c$. Introduce fields $j_{t}=\partial_{t}\varphi$ and $j_{x}=\partial_{x}\varphi$. As the dependence of $S[\varphi]$ on $\varphi$ only enters through $j$, we would like to replace the path integral over $\varphi$ with a path integral over $j$; however, in doing so we need to restrict to field configurations satisfying $\partial_{t}j_{x}=\partial_{x}j_{t}$. This is accomplished with a Lagrange multiplier $\theta$, leading to the action
$S[j,\theta]=\int\mathrm{d}x\mathrm{d}t~{}\frac{R^{2}}{4\pi}(j_{t}^{2}-j_{x}^{2})+\frac{\theta}{2\pi}(\partial_{t}j_{x}-\partial_{x}j_{t}).$ (92)
Compactness of $\theta\sim\theta+2\pi$ ensures that vortex configurations of $j$ are included in the path integral. Integrating out $j$ gives the action in terms of the dual $\theta$
$S[\theta]=\frac{1}{4\pi R^{2}}\int\mathrm{d}x\mathrm{d}t~{}(\partial_{t}\theta)^{2}-(\partial_{x}\theta)^{2}$ (93)
where the symmetry $\theta\to\theta+c$ is manifest. This action, like (91), does not have a reparametrization invariance.
Let us return to (92) and consider the modified action
$S[\varphi,j_{x},\theta]=\int\mathrm{d}x\mathrm{d}t~{}\frac{R^{2}}{4\pi}((\partial_{t}\varphi)^{2}-j_{x}^{2})+\frac{\theta}{2\pi}(\partial_{t}j_{x}-\partial_{x}\partial_{t}\varphi)$ (94)
where we have replaced $j_{t}$ with $\partial_{t}\theta$ but kept $j_{x}$ in anticipation of integrating it out. Interestingly, this action has two reparametrization symmetries
$\varphi\to\varphi+f(x),~{}\theta\to\theta+g(x)$ (95)
for time-independent functions $f(x)$ and $g(x)$. Integrating out $j_{x}$, we obtain
$S[\varphi,\theta]=\int\mathrm{d}x\mathrm{d}t~{}\frac{1}{2\pi}\partial_{t}\varphi\partial_{x}\theta+\frac{R^{2}}{4\pi}(\partial_{t}\varphi)^{2}+\frac{1}{4\pi R^{2}}(\partial_{t}\theta)^{2}.$ (96)
This can be thought of implementing “half” of T-duality to rewrite the action in terms of both $\varphi$ and $\theta$ variables in such a way that both $\varphi$ and $\theta$ have a shift symmetry. This action is in standard “hydrodynamic form”: the first term characterizes the mixed anomaly between the two $U(1)$ symmetries, and the remaining terms depend only on time derivatives of the phase fields, which is standard when the symmetries are unbroken. This occurred naturally when both $U(1)$ symmetries were made manifest in the action. Note that the reparametrization symmetry is not accidental: relaxing the symmetry and allowing terms such as $(\partial_{x}\theta)^{2}$ changes the action and makes the theory different from the compact boson and Dirac fermion.
Some care is required in going between this action and the original compact boson action that we started with. Varying with respect to $\theta$ of (96) gives
$\partial_{t}\left(\partial_{x}\varphi+\frac{1}{R^{2}}\partial_{t}\theta\right)=0$ (97)
which is solved by setting $\partial_{x}\varphi+\frac{1}{R^{2}}\partial_{t}\theta=F(x)$. Using the equations of motion to integrate out $\theta$ and shifting $S\to S-\int\mathrm{d}x\mathrm{d}t~{}\frac{1}{2\pi}F(x)\partial_{t}\theta$ by a total derivative term gives
$S[\varphi]=\frac{R^{2}}{4\pi}\int\mathrm{d}x\mathrm{d}t~{}(\partial_{t}\varphi)^{2}-(\partial_{x}\varphi-F)^{2}$ (98)
which reduces to the compact boson action when $F=0$. Now, $S[\varphi]$ can be invariant under the reparameterization symmetry $\varphi\to\varphi+f(x)$ if we also shift $F\to F-\partial_{x}f(x)$. This suggests that different initial spatial profiles of $\varphi$ do not change the physics of Dirac fermions; indeed, one can add a constant phase to the textbook bosonization dictionary Shankar (2017) without causing any change to the correlation function and the commutation relation. By contrast, this initial condition will cost energy in superfluids where the reparameterization symmetry is absent.